F(x)=(2/3)^x
Translate down 5 units
Also
Translate up 2 units left 4 units

Answer by stanbon(75887)

F(x)=(2/3)^x
Translate down 5 units
G(x) = (2/3)^x -5
---------------------------
Also
Translate up 2 units left 4 units
H(x) = (2/3)^(x+4) +2
=============
cheers,
Stan H.